V.L. Barnes and Company, a manufacturer of quality handmade walnut bowls, has had a steady growth in sales for the past 5 yeai Data Table Variable cost (per bowl) Direct materials Direct manufacturing labor $ Variable overhead (manufacturing, marketiI just need the answers for 3-6 please! thank you :)
Answer #1

Problem 3

Particulars Amount units price per unit Total
sales 24000 33 792000
less ; variable cost 24000 16.5 396000
Direct Materials 3.5
Direct Manufacturing labour 8.5
variable overhead 4.5
contribution 396000
less ; Fixed cost
Manufacturing 18000
marketing & distribution 196500
Additional marketing 29700
total Fixed Cost 244200
Net operating income 151800
less ; income tax (@40% on net operating income 60720
Net Income 91080

problem 6

maximum amount that can be spent on advertising to earn profit of 90330

net operating income = 90330/60% = 150550

maximum amt spent on advertising in addition to the 29700 = 151800 - 150550 = 1250

total additional expense that can be sent = 29700+1250 = 30950

problemy In 2018 fixed cost = 244200 contribution per unit = 33-16.5 216.5 cm ratio: 16.5 100- 50% YEPE Qu4zo0 Bolak even (in

Problem 5

Marginal Income statement for year 2017

Particulars

Amount

units

price per unit

Total

sales

23000

33

759000

less ;

variable cost

23000

16.5

379500

Direct Materials

3.5

Direct Manufacturing labour

8.5

variable overhead

4.5

contribution

379500

less ;

Fixed cost

Manufacturing

18000

marketing & distribution

196500

total Fixed Cost

214500

Net operating income

165000

less ;

income tax (@40% on net operating income

66000

Net Income

99000

To obtain same income as per 2017

           Net operating income in 2018 = 165000

Required revenue = ( fixed cost in 2018 + profit ) / cm ratio in 2018

                               = ( 244200+165000) / 50%

                             =        818400